A store mixes bluegrass worth $12 per pound and ryegrass $16 per pound. The mixture is to sell for $13 per pound. Find out how much of each should be used to make a 558 pound mixture?
Answer by addingup(3677)

bluegrass: x
ryegrass: y
x+y = 558 subtract x from both sides:
y = 558-x
------------------------------
12x+16y = 13(558) substitute for y:
12x+16(558-x) = 13(558)
12x+8928-16x = 7254
-4x = -1674
x = -1674/-4 divide, and remember that -/- = +
x = 418.50
You need 418.5 pounds of the bluegrass and 558-418.5= 139.5 pounds of the ryegrass
------------------------------------------
make-sense check:
Mixture price: 13
Look at the bluegrass: 12, very close to 13
Look at the rye grass: 16, not so close to 13
It makes sense that you need a lot more of the $12 bluegrass
:
Happy learning,
John